OpenStudy (anonymous):

Can someone help me with this problem? I don't even know what it's called to look it up on a math website. 5x^3-2 if x<-6

OpenStudy (anonymous):

umm do u want it graphed

OpenStudy (anonymous):

or solved?

OpenStudy (anonymous):

cause i dont know what part to solve

OpenStudy (anonymous):

This is a math packet I'm supposed to do as an incoming junior in a precalculus class. The only instructions are to how al my work on a separate sheet of paper and consult purplemath.com if help is needed

OpenStudy (anonymous):

i know the graph 5(x^3)-2 where u restrict the domain to x<-6

OpenStudy (dumbcow):

well you could substitute in -6 to give a limit to the function -> 5*(-6)^3 -2 = -1082 so f(x) < -1082 for x<-6

OpenStudy (anonymous):

yh can restrict y values with it as well

OpenStudy (dumbcow):

i recommend graphing it

OpenStudy (anonymous):

the answer sheet says the answer is 106

OpenStudy (anonymous):

but here's the whole problem: let f(x) = [ 5x^3 -2 if x <-6 3x^2 - 2 if -6 <_ x <7 7+ 2x if x _> 7 find f(-6), f(0), f(6), f(7), f(10)

OpenStudy (anonymous):

http://www.purplemath.com/modules/ineqsolv.htm

OpenStudy (anonymous):

its is a piece wise function

OpenStudy (anonymous):

u plug in f(-6) in second equation 3(-6)^2-2 because x is between -6 and 7=106 u plug in f(0) in second equation 3(0)^2-2 because x is between -6 and 7=0 u plug in f(7) in third equation 7+2(7) because x is = 7 u plug in f(10) in third equation 7+2(10) because x is more than 7

OpenStudy (anonymous):

answers are 106

OpenStudy (anonymous):

-2

OpenStudy (anonymous):

21

OpenStudy (anonymous):

27

OpenStudy (anonymous):

THANK YOU SO MUCH!!!!